OpenStudy (anonymous):

find the derivative of y = e^(k tan(7x)^(1/2))

OpenStudy (anonymous):

\[\huge e^{ktan(7x)^{\frac{1}{2}}}?\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\int\limits \frac{7 k \sec ^2(7 x) e^{k \sqrt{\tan (7 x)}}}{2 \sqrt{\tan (7 x)}} \, dx=e^{k \sqrt{\tan (7 x)}} \]

OpenStudy (anonymous):

...the question is \[e ^{\tan \sqrt{7x}}\]

OpenStudy (anonymous):

The thing on the left (minus the integral which is this \[\int\limits_{}^{}\] Is the answer

OpenStudy (anonymous):

\[\Large \frac{7 k \sec ^2(7 x) e^{k \sqrt{\tan (7 x)}}}{2 \sqrt{\tan (7 x)}}\]